A body of mass 6kg is at rest when a force of magnitude 30 N on the body. After 10s what will be kinetic energy?
1
Expert's answer
2019-02-04T10:11:08-0500

Find the acceleration of the body using Newton's second law

a=F/m=30/6=5 m/s^2 .

According to the condition of the problem initial body speed v_0=0 m/v. Equation of speed with uniformly accelerated motion

v=v_0+at,

where t is time. Therefore, after 10s speed of the body is equal to

v=0+5∙10=50 m/s

Hence kinetic energy after 10s is equal to

E=(mv^2)/2=(6∙〖50〗^2)/2 7500J

where m=6kg is mass of the body, v is speed.